OpenStudy (anonymous):

suppose you want to earn $80. What is the difference of the maximum number of hours and minimnum number of hours you need to work to make $80 if you are paid at two different rates?

OpenStudy (bahrom7893):

well what are the rates?

OpenStudy (bahrom7893):

There's a difference when I get paid $0.01 per hour vs $1 per hour vs $10 per hour

OpenStudy (anonymous):

Well, it would help if we have one of the rates

OpenStudy (anonymous):

Oh right sorry. rate a=$5/h rate b=$8/hr

OpenStudy (bahrom7893):

80/5 = x hours 80/8 = y hours The difference is: x-y hours

OpenStudy (anonymous):

so it would be 10 hours? :)

OpenStudy (bahrom7893):

(80/5)-(80/8) = 16 - 10 How did you get 10 hours?

OpenStudy (anonymous):

Oh... I just took your 80/8 and divided it and got 8.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

oh nvm..its 6 hours lol

OpenStudy (anonymous):

thanks for your help!

OpenStudy (anonymous):

Your welcome for trying to help

OpenStudy (anonymous):

:)